OpenStudy (anonymous):

1. TNT, (C7H5N3O6) reacts explosively with oxygen gas (O2) to produce CO2 gas, water vapor, and N2 gas. If James weighs out 0.853 g of TNT, answer the following question: How many mg of O2 will react with the TNT?

OpenStudy (abb0t):

TNT (also known as trinitrotoluene, or 2-Methyl-1,3,5-trinitrobenzene) has the molecular formula: C\(_7\)H\(_5\)N\(_3\)O\(_6\). And hence, when reacted in oxygen gas, you get what is known as \(combustion\) reaction. the reaction is: \(\sf \color{red}{C_7H_5N_3O_6 + O_2 \rightarrow CO_2 + N_2 + H_2O} \)

OpenStudy (anonymous):

I see, that is what I thought...my text has very few examples of finding the mg of O2 that will react or the amount of H2O produced. I balanced the reaction to look like 4C7H5N3O6 + 21O2 = 28CO2 + 10H2O + 6N2 but now I'm stuck!

OpenStudy (abb0t):

So, you start always with what you are given, meaning \(\sf \color{red}{0.853~g~TNT}\) and go from there.

OpenStudy (anonymous):

I found the molar mass of the TNT to be 227.14g/mol

OpenStudy (abb0t):

But before you even begin making conversions, make sure that your chemical equation is BALANCED!

OpenStudy (anonymous):

So, when I am calculating the molar mass to do conversions I have to use the balanced equation...I think that is where I'm getting confused. Do I multiply the molar mass by the coefficient i get as the result of balancing?

OpenStudy (abb0t):

Yes. You first have have your equation balanced so that you know the ratio that will be used to convert from 1 to the other.

OpenStudy (abb0t):

By the way, this is such an unusual example, a student would never be working with TNT in a laboratory. Lol. It's highly reactive.

OpenStudy (anonymous):

These are my lab exercises...the Professor doing the lectures is not giving us examples like these so I have little to work with. Once I balance the equation I get 21O2, I use that converted into g/Mol then into mg?

OpenStudy (anonymous):

Thanks, sorry got disconnected. Trying to solve this now...

OpenStudy (abb0t):

Yeah. So \(\sf \color{red}{0.953~g~TNT} \times \frac{1~mol~TNT}{mass~of~TNT} \times \frac{number~mol~O_2}{number~mol~of~TNT} \times \frac{grams~of~O_2}{1~mol~O_2} \times \frac{1000~mg}{1~g}\) = _______ mg of O\(_2\)